OpenStudy (anonymous):

Find the exact value of sin(T -pi/6) when sin(t)=1/3 and cos(t)=Rad8/3

OpenStudy (amistre64):

\[sin(a+b)=sin(a)cos(b)+sin(b)cos(a)\]

OpenStudy (amistre64):

in this case: a = T and b = -pi/6 but not sure if that fits too well

OpenStudy (amistre64):

that might work, even with the funny looking decimal

OpenStudy (anonymous):

ops there is a error there i will correct that

myininaya (myininaya):

what funny looking decimal?

myininaya (myininaya):

T=t?

OpenStudy (anonymous):

\[\sin(T - \pi/6) = \sin(T)\cos(\pi/6)-\sin(\pi/6)\cos(T)\]\[\sqrt{3}/6 - 0,0232 = 0,2654\]

OpenStudy (anonymous):

so basically use the double angle formula?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[\frac{1}{3} \cdot \frac{\sqrt{3}}{2}-\frac{1}{2} \cdot \frac{\sqrt{8}}{3}\]

OpenStudy (anonymous):

oh did you get that from a chart myininaya?

OpenStudy (anonymous):

Remember that \[\sin(a \pm b)= \sin(a)\cos(b) \pm \sin(b)\cos(a)\]

OpenStudy (anonymous):

ok.. yes

myininaya (myininaya):

yep bella \[\cos(\frac{\pi}{6})=\frac{\sqrt{3}}{2}; \sin(\frac{\pi}{6})=\frac{1}{2}\]

OpenStudy (anonymous):

ok thank you!

OpenStudy (amistre64):

across the pond they use the "," as a "."